Circular No. 5680

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET OHSHITA (1992a_{1})

Ephemeris continuation from IAUC 5679:

1993TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Jan. 3 15^{h}14.^{m}40 +60^{o}35.^{′}4 0.878 1.369 94.^{o}6 45.^{o}7 13.1
5 15 30.89 +63 45.0 0.901 1.399 95.8 44.4 13.2
7 15 49.87 +66 38.5 0.927 1.428 96.8 43.1 13.4
9 16 11.79 +69 14.3 0.956 1.458 97.5 42.0 13.5
11 16 37.1 +71 30.8 0.988 1.487 97.9 40.9 13.7
13 17 06.1 +73 26.6 1.023 1.517 98.2 39.9 13.9
15 17 38.8 +75 00.5 1.060 1.546 98.3 39.0 14.0
17 18 14.9 +76 11.6 1.098 1.576 98.2 38.2 14.2
19 18 53.2 +76 59.9 1.139 1.605 97.9 37.4 14.3
21 19 32.0 +77 26.7 1.181 1.634 97.6 36.7 14.5
23 20 09.6 +77 34.4 1.225 1.663 97.1 36.0 14.6
25 20 44.5 +77 26.7 1.270 1.692 96.5 35.3 14.8
27 21 15.9 +77 07.3 1.316 1.721 95.8 34.7 15.0
29 21 43.6 +76 39.6 1.363 1.750 95.0 34.1 15.1
31 22 07.7 +76 06.7 1.410 1.778 94.2 33.5 15.2
Feb. 2 22 28.6 +75 30.8 1.459 1.807 93.3 33.0 15.4

Total V CCD magnitude obtained by H. Mikuz, Ljubljana, Slovenia: Dec. 7.19 UT, 11.9 (2′.8 coma).

EY CYGNI M. Orio, Torino Observatory; and H. Ogelman, University of Wisconsin

at Madison, communicate: “We report the detection of the dwarf nova EY Cyg in soft x-rays with the Rosat PSPC during a serendipitous pointing on 1991 Apr. 5. The x-ray source was ∼ 0^{o}.6 off-axis and had a net vignetting corrected countrate of 0.08 ± 0.01 cts/s, corresponding to an energy flux at the detector of ∼ 3 × 10^{−13} erg cm^{−2} s^{−1} in the energy interval 0.4-2.2 keV. This is likely to be the flux at quiescence, since no outburst was reported at the epoch of the detection.”

PERIODIC COMET GICLAS (1992l) Total visual magnitude estimates by A. Nakamura, Kuma, Japan (0.60-

m reflector): Nov. 30.69 UT, 14.9; Dec. 2.68, 14.8.

1992 December 21 Daniel W. E. Green